OpenStudy (anonymous):

The sum of two consecutive odd numbers is 20. What are the two numbers? Which let statement would you use?

OpenStudy (anonymous):

HEY COULD YALL HELP ME ON fashiongirl08092002's question 

OpenStudy (anonymous):

x + (x+2) = 20 x + x + 2 = 20 (x + x = 2x) + 2 = 20 2x + 2 = 20 2x + 2 - 2 = 20 - 2 2x = 18 2x/2 = 18/2 x = 9 The value of x is 9 so the first number is 9, the second is 11. P.S. Formula for finding 2 ODD consecutive numbers are x + (x+2)

OpenStudy (anonymous):

Let x represent the number(s).